abbbbbb:

byron is 3 years older than doug the product of their ages is 40 how old is doug

darkknight:

So lets say bryon is b and doug is d So we know that bd=40 and b=d+3 2 variables requires 2 equations substitue b for d+3 d+3)(d)=40 d^2+3d=40 d^2+3d-40=0 d-5)(d+8)=0 d=5 or -8 age can't be negative so d=5. and b is 5+3 or 8

ramen:

or you can literally just find numbers that multiply to 40 but have a difference of 3. so 8-5 = 3 and 8*5=40